OpenStudy (anonymous):

Find the equation for the line perpindicular to 7x-21y=39 that passes through the point (-3, 5). answer in y=mx+b. Simplify as much as possible, but leave iproper fractiobns. help please?

OpenStudy (anonymous):

21y = 7x - 39 y = (1/3) x - 13/3 slope of the perpendicular line = -1 / (1/3) = -3 using y-y1 = m(x-x1) y - 5 = -3(x+3) y = -3x -4

OpenStudy (anonymous):

alright thank you again, sure it won't be the last